During its first month of operations, Purrfect Pets purchased 5,800 bags of dog food at a cost of $2 a bag and sold all 5,800 bags of dog food on account with payment terms of $ 1/10, net 30 for $10 each. A total of 2,800 of these bags were sold to customers who paid within the discount period; the other customers paid after the discount period had ended. Sales allowances totaling $200 were granted to customers whose dogs did not like the dog food a. Calculate the gross profit for the month. b. Calculate the gross profit percentage for the month. (Round your answer to one decimal place (i.e., 0.1234 should be entered as 12.3)).
Answer #1

Cost of Bags sold = 5800 * 2 = $11600

Discount allowed on sales = 2800 *10 *0.01 = 280

sales = 5800 *10= $58000

A) Gross profit = Net sales -COGS

              = (58000 - 280 -200) - 11600

              = $45,920

b)Gross profit % = Gross profit /net sales

                          = $45,920 / $58000

                         = 79.17%